### v1

Abstract
Representations from large language models are highly effective at predicting BOLD fMRI responses to language stimuli. However, these representations are largely opaque: it is unclear what features of the language stimulus drive the response in each brain area. We present generative explanation-mediated validation (GEM-V), a framework for generating concise explanations of language selectivity in the brain and then validating those explanations in follow-up experiments that use synthetic stimuli. This approach is successful at explaining selectivity both in individual voxels and cortical regions of interest (ROIs). We show that explanatory accuracy is closely related to the predictive power and stability of the underlying statistical models. These results demonstrate that LLMs can be used to bridge the widening gap between data-driven models and formal scientific theories.
keywords:
Language models, Encoding models, fMRI, Synthetic data
1   Introduction
Science faces an explainability crisis: data-driven deep learning methods are proving capable of modeling many natural phenomena, like protein folding, meteorological events, and computations in the brain [1, 2, 3]. However, these models are not scientific theories that describe the world in natural language. Instead, they are implemented in the form of vast neural networks with millions or billions of largely inscrutable parameters. One emblematic field is language neuroscience, where large language models (LLMs) are highly effective at predicting human brain responses to natural language, but are virtually impossible to interpret or analyze by hand [4, 5, 6, 7, 8, 9, 10].
To overcome this challenge, we introduce the generative explanation-mediated validation (GEM-V) framework. GEM-V translates deep learning models of language selectivity in the brain into concise verbal explanations, and then designs follow-up experiments to verify that these explanations are causally related to brain activity.
Figure 1: Driving single-voxel response with generative explanation-mediated validation. (a) Voxelwise BOLD responses were recorded using fMRI as human subjects listened to 20 hours of narrative stories. An encoding model $f$ was fit to predict these responses from the story text. $f$ consists of a linear model fit on representations extracted from an LLM, which are not readily interpretable. Encoding models were tested by predicting responses on held-out fMRI data, and only well-performing models were selected for further analysis. (b) We used an automated procedure to find a verbal description of the function that $f$ computes for each voxel. First, we tested $f$ on a large catalog of $n$-grams ($n=1,2,3)$ and found those that maximally drove predicted responses. These $n$-grams were then summarized into stable explanation candidates using a powerful instruction-tuned LLM. Finally, we evaluated each explanation candidate by generating corresponding synthetic sentences and testing that these sentences yielded large predictions from $f$. (c) To test whether the generated explanations were causally related to activation in the brain, we used an LLM to produce synthetic narrative stories where each paragraph is designed to drive responses based on the generated explanation for one voxel. For each subject we constructed stories to drive 17 well-modeled voxels with diverse selectivity. These stories were then presented to the subjects in a second fMRI experiment. (d) Average BOLD response during its driving paragraph for each voxel, relative to baseline, i.e. average response to all non-driving paragraphs. On average, driven responses were significantly higher than baseline for each subject ($p=0.020$ (S01), $p<10^{-5}$ (S02), $p=0.009$ (S03); permutation test, FDR-corrected). For well-driven voxels, this means that the generated explanation is causally related to activation of that voxel, and thus that we have successfully translated the LLM-based encoding model into a verbal explanation. (e) Average BOLD response for each selected voxel to each of the driving paragraphs in one subject (S02). Responses to the driving paragraph generated using the explanation for that voxel appear along the main diagonal. Explanations that were used to construct the driving paragraphs are shown below. BOLD responses were generally high for the driving paragraphs for each voxel as well as semantically related paragraphs (e.g. directions and locations, emotional expression and laughter).
2   Language encoding models
We designed GEM-V to interpret language encoding models, which are data-driven, LLM-based models of cortical language selectivity. We thus began by fitting encoding models (Fig. 1a) for each voxel in each of 3 human subjects using 20 hours of passive language listening fMRI data per subject collected in a prior study [11]. The 20 hours of speech stimuli were transcribed and then passed through multiple open-source large language models [12, 13] to obtain activation vectors for each word. Multiple models were built using different LLMs to ensure that predictions are stable [14, 15], i.e. not idiosyncratic to a single model. Regularized linear regression was then used to predict the response timecourse of each voxel as a weighted combination of LLM activations. Encoding models were tested by predicting responses on held out fMRI data and then computing the correlation between predicted and actual responses. Encoding models using advanced LLMs achieve very high prediction performance [8]; here, only models with sufficient prediction performance ($r>0.15$) were used in further analyses, with many voxels predicted at $r>0.7$. Full preprocessing, modeling, and data collection details are provided in the Methods.
3   Generating and evaluating natural-language explanations for single voxels
Each encoding model represents the language selectivity of one part of the brain, but does so as a linear combination of LLM activations that are not human-interpretable. The first step in GEM-V is to convert these models into concise natural language explanations: a word or short phrase summarizing the properties of language that drive responses most strongly. This was accomplished by using an instruction-finetuned LLM [16] to generate candidate explanations by summarizing the n-grams that yield the largest predictions from the model (Fig. 1b) [17]. The resulting explanation is given in easily understandable natural language. For example, we found that some voxels appeared to be selective for language describing Food preparation (Fig. 1b). To ensure that the generated explanation correctly captures the function computed by the encoding model, we used the same LLM to generate synthetic text from the explanation and checked that this text could successfully drive the voxel encoding model. However, this does not guarantee that this function is correctly aligned to the activity of the brain in the real world.
To fully close the loop and confirm that a generated explanation accurately reflects language selectivity in the brain, the second step of GEM-V is to automatically design a new neuroimaging experiment to test the generated explanation in vivo. This was done by prompting an instruction-finetuned LLM [16] to generate narratives that should selectively drive cortical activation based on that explanation (Fig. 1c). If the explanation generated for a voxel is accurate, then text generated according to that explanation should elicit large responses in that voxel; this would demonstrate that the explanation has causal influence over that voxel’s activity. Given a selected set of voxels and their explanations, we constructed narrative stories by iteratively prompting an LLM [16] to prioritize a different voxel’s explanation as the main focus of each paragraph. This allowed us to test whether voxels selectively respond to paragraphs that match their explanation. The LLM ensures these stories remain coherent and engaging, helping to keep subjects attentive during the fMRI experiment [18, 7].
We used GEM-V to generate and test explanations for 17 voxels with strong encoding model predictive performance in each of 3 subjects. The number of voxels was selected so that resulting narratives would be similar in length to the stimuli used to initially fit the encoding models [11]. We then measured the average response of each target voxel during the paragraph that was designed to drive it and compared this to the average response to other paragraphs (Fig. 1d). Across voxels, responses to driving paragraphs were significantly greater than baseline responses for each subject ($p=0.020$ (S01), $p<10^{-5}$ (S02), $p=0.009$ (S03); permutation test with Benjamini-Hochberg false discovery rate correction [19]). Of the 51 tested voxels across the 3 subjects, 41 show increased response, with an average increase of 0.198 standard deviations over baseline. Differences in effect size are related to head motion of subjects during test time and variation in sample size (Table 2). This demonstrates that the generated explanations are mostly effective drivers of brain activity for their chosen voxels. Further, because this experiment was not based directly on the LLM encoding model [9] but was designed using the generated explanation, this suggests that we have successfully found explanations that are causally linked to brain activity.
The results in Fig. 1d compare responses during the driving paragraphs generated for each voxel to average responses over all other paragraphs. But if two voxels have very similar explanations and both are correct, then we would expect each to also be driven by the other’s generated stimuli. To explore this question we disaggregated the results by showing the response of each voxel to every driving paragraph in one subject (Fig. 1e). Most voxels are driven by their own explanations, as can be seen by the positive values on the main diagonal. However, many voxels are also strongly driven by related explanations (e.g. directions, measurements, and locations; communication and emotional expression). This demonstrates that, in most cases, semantically related explanations will drive similar sets of voxels. We find that the same setting succeeds in driving pairs of voxels rather than individual voxels (Section A.5).
Figure 2: Driving ROI response with generative explanation-mediated validation for subject S02. (a) Explanations were generated and used to drive 8 well-defined regions of interest. Responses in all ROIs were significantly driven above baseline ($p<0.05$; permutation test, FDR-corrected). (b) To understand driving performance with more granularity, we color each voxel in each ROI by how well it was driven by its corresponding driving paragraph. The resulting composite flatmap occasionally shows subregions within ROIs that are more selectively driven for a particular explanation. (c) GEM-V can also be used to build more nuanced theories of cortical semantic selectivity. We focused on three ROIs that are known to have similar selectivity for place concepts: retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). When explanations were generated for each ROI independently we found that each ROI was driven by all three driving paragraphs (left side). To distinguish these ROIs, we used GEM-V to find new explanations and construct stories that would selectively drive each area while suppressing the other two. Testing these stories in an fMRI experiment showed that we succeeded in finding selective explanations for two ROIs: RSC is selectively driven by location names and PPA by unappetizing foods. However, the explanation for OPA, spatial positioning & directions, drove responses in all three ROIs (right side). (d) Visualization of the place area driving experiment with voxel-level granularity. We show a 3-channel flatmap showing the outcome of each location-selective driving experiment; a voxel is more red/green/blue if that voxel was driven by the corresponding ROI explanation.
4   Explaining selectivity in regions of interest (ROIs)
Theories of language selectivity often revolve around regions of interest (ROIs) that are believed to be selective for specific semantic categories [20], such as locations (OPA) [21, 22] or body parts (EBA) [23, 24]. We explored whether GEM-V could be used to independently uncover the semantic selectivity of regions of interest, both for generating new hypotheses as well as re-establishing known functional selectivity. We used GEM-V to find explanations and design synthetic driving stimuli for a diverse selection of individual ROIs that were localized using separate stimuli: the extrastriate body area (EBA), intraparietal sulcus (IPS), occipital face area (OFA), posterior superior temporal sulcus (pSTS), superior premotor ventral (sPMv), retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). These ROIs were identified in each subject using separate localizer scans. Fig. 2a shows the generated explanations and average response above baseline for each ROI during each driving experiment. The explanations found by GEM-V and validated in the follow-up experiment are well-matched to known selectivity, e.g. Body parts in EBA, and Scenes and settings in PPA. Driving succeeded for every ROI (all $p<0.05$; permutation test with FDR correction), although for some ROIs, e.g. RSC, the effect is stronger. Fig. 2b visualizes driving responses for each ROI on a composite flatmap, where each voxel is colored on the basis of how much it was driven by its ROI’s explanation. This map suggests that the variability in driving between ROIs reflects functional heterogeneity within each region.
We next tested where GEM-V could be used to describe differences between regions that appear to have similar selectivity [20]. Examining three ROIs known to be selective for places or locations (RSC, PPA, OPA) [25], we found that the driving paragraphs designed independently for each ROI also drove responses in the others (Fig. 2c left). To differentiate these ROIs, we used GEM-V to generate new explanations that should selectively drive each one of these regions while suppressing the other two [26]. These selective explanations deviated somewhat from the original ROI explanations, e.g. for RSC the explanation changed from Travel and location names to just Location names, suggesting that the travel part of the explanation is common across ROIs while location names are more unique to RSC. New stories were generated to include only instances of the corresponding explanation, for instance, when driving only RSC, location names were included, while directions were specifically excluded (see prompts in Section A.1). Since these two categories commonly cooccur in natural language, these synthetic stimuli enabled us to examine differences that might not be present in typical naturalistic studies. Testing these stories in another fMRI experiment showed that GEM-V was able to find explanations that can selectively drive two of the ROIs—Location names for RSC and Unappetizing foods for PPA—while the explanation of Spatial positioning & directions for OPA still drove responses in all three ROIs (Fig. 2c right; Fig. 2d). These results demonstrate that GEM-V can be used to build more nuanced theories of cortical semantic selectivity beyond general trends that are well-understood.
Figure 3: Responses to driving paragraphs reproduce known semantic contrasts and open new hypotheses. To understand how each explanation activates the entire cortex, we visualize the driven response (average normalized BOLD response) for six explanations. (a) Some maps recovered well-established results, e.g. the explanation Locations gave strong responses in the place areas RSC, OPA, and PPA. (b) Other maps independently confirmed newer hypotheses, e.g. the explanation Food Preparation activates a region in ventral occipital cortex near the fusiform face area (FFA) [27, 28]. (c) Finally, some explanations did not clearly map to any known hypotheses, but may suggest new directions for future research. All observed activation patterns along with ROI localization explanations are presented in Appendix A.7.
5   Responses to synthetic stories reveal underlying contrast effects
Although the stories generated by GEM-V are designed to activate single voxels or ROIs, we can measure how much each explanation activates every part of cortex. These experiments can thus be seen as a hybrid between classical block-designs that are easy to analyze and natural stimulus experiments that elicit stronger and more widespread activation in cortex [18, 7]. To demonstrate this, we computed the average activation map for each explanation by averaging responses over timepoints within each paragraph. The results for six selected explanations in one subject are shown in Fig. 3.
Some explanations were similar to previously well-established contrasts (Fig. 3a) and gave the expected results. For example, the explanation Locations gave strong responses in the place areas RSC, OPA, and PPA [25]. Others confirmed newer hypotheses. For example, the Food Preparation explanation (Fig. 3b Top) activates a region in ventral occipital cortex near the fusiform face area (FFA). Selectivity for this area to food was found recently in responses to images [27, 28], but these results demonstrate for the first time that this selectivity also extends to linguistic descriptions of food. Finally, some explanations did not clearly map to any known hypotheses (Fig. 3c), but still demonstrate suggestive patterns of activation that could be explored in future experiments. Maps for other explanations are provided in Appendix A.7.
6   Factors underlying effective GEM-V stories and voxels
Figure 4: Analyzing factors that impact explanation-mediated validation. To evaluate whether GEM-V succeeds in generating effective stimulus stories, we assessed the driving paragraphs of the stimulus. (a) To confirm that generated paragraphs match the explanation used to construct them, a matching score was computed for each explanation and paragraph by using an LLM to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation and then z-scoring the result. Each driving paragraph showed a strong match with its generating explanation. Plot shows one subject (S02), similar plots for other subjects are shown in Section A.4. (b) To confirm that each driving paragraph effectively drives its corresponding encoding model predictive performance, we computed the predicted response in each selected voxel to each generated paragraph. This revealed strong matches for most voxels, along with some matches between driving paragraphs and voxels with semantically similar explanations, e.g. directions and locations. Plot shows one subject (S02), similar plots for other subjects are shown in Section A.4. (c) After running the fMRI driving experiment, we found that a key factor determining whether a voxel is driven well by the GEM-V framework was the stability score for the voxel, i.e. the correlation between the n-gram rankings provided by the LLaMA-based encoding model and the OPT-based encoding model. (d) Another important factor for eliciting increased driving responses is the presence of key n-grams in the driving paragraphs. These n-grams induce a standard hemodynamic response curve that peaks at around 6 seconds, yielding a significant increase ($p=0.009$; one-sided t-test). (e) Finally, to test whether the driving results were sensitive to the particular voxels that were selected, we evaluated whether the GEM-V stories drove alternative voxels in each subject that were assigned the same explanation as the target voxels being driven. Both the targed voxels and alternative voxels also showed significantly increased driving responses ($p<0.05$; permutation test, FDR-corrected).
The experiments thus far showed that GEM-V can effectively explain and drive a variety of locations throughout cortex, but fails to successfully drive some voxels. To understand these limitations, we evaluated factors that influence driving performance, beginning with factors surrounding the generation of story stimuli.
First, we considered whether driving failures may arise from a semantic mismatch between the proposed explanation and the generated story stimulus. We measured the match between each driving paragraph and its target explanation for subject S02 by using an LLM to evaluate similarity between the explanation and trigrams in the driving paragraph (see Methods). There is a strong correspondence between each voxel explanation and its driving paragraph (Fig. 4a, orange diagonal), suggesting that LLM stimulus generation does not account for the failures. Moreover, there is no clear correlation between these scores and driving success (average correlation of -0.05; see Fig. 9a). Second, we tested whether the failed cases result from a misalignment between the original encoding model for each voxel and the generated story stimulus. Again, each voxel’s encoding model showed an increased response for its driving paragraph (Fig. 4b, orange diagonal), suggesting the driving paragraphs are successfully aligned with the encoding model. In this case, higher driving scores for the encoding model did yield higher driving scores in the follow-up experiment for 2 of the 3 subjects (average correlation of 0.19; see Fig. 9b).
Given that driving failures are neither the result of stimulus-explanation mismatches nor stimulus-model mismatches, we concluded that they must largely stem from limitations in the encoding models. We only applied GEM-V to voxels with high encoding model test performance, but this does not guarantee that the encoding model is accurate. Model performance is measured on a dataset that is limited in scope and that comprises real, noisy data. Both of these factors could cause a voxel to appear well-predicted despite having an inaccurate encoding model, or to appear poorly-predicted despite having an accurate encoding model. Since these limitations cannot be addressed directly, we investigated factors that improve confidence in the generated explanations by assessing their stability, a key principle underlying effective statistical interpretation according to the predictability, computability, and stability (PCS) framework [14, 29].
To identify which explanations are stable, we defined a stability score that measures agreement between the predictions of two encoding models built from different LLMs (see Methods) for the same voxel using the same fMRI data. We found a strong positive correlation between the stability score of a voxel and its mean driving response across all three subjects (Fig. 4c), suggesting that the stability score reliably informs the causal reliability of an explanation. This result underscores the idea that models must be predictive and stable in order to effectively generate hypotheses for follow-up experiments [15, 30].
We also assessed the stability of our framework to many modeling choices including the prompts used to generate synthetic stories (Fig. 7) and the granularity of units we drive (e.g. single voxels versus semantic clusters of voxels in Fig. 8). In both cases, we found that the magnitude of the evoked driving responses may change, but the overall trends in driving voxels are consistent.
To further understand how driving paragraphs succeed, we measured whether the inclusion of key driving n-grams extracted from the encoding model drove responses in a temporally specific fashion (Fig. 1d). During the fMRI experiment, the presentation of these n-grams evoked a significant increase in responses, peaking 6 seconds after presentation (Fig. 4d; $p=0.009$; one-sided t-test). This strongly aligns with hemodynamic response curves which tend to peak about six seconds after stimulus presentation [31], and are further validation that the generated stimuli are the cause of the driving effects we observe.
[…]
7   Discussion
Many prior studies have demonstrated the predictive power of black-box encoding models in language neuroscience across modalities such as fMRI, ECoG, EEG, and MEG [32, 33, 4, 5, 8], but these models have been criticized for their inability to produce accurate and reliable explanations [34] and limited efforts have been made to explicitly interpret these models [35, 36, 37]. Our work takes the first steps in overcoming this criticism by enabling us to directly test the causality of an encoding model explanation through the use of LLM-generated synthetic stimuli.
This work is related to works that have generated stimuli to drive neurons in the visual cortex [38, 15, 39], as well as more recent work that has used decoding to provide insight into visual cortex selectivity [40]. In language fMRI, the closest work is Tuckute et al. 2023, which demonstrated the ability to drive the language network as a whole (rather than for individual voxels or ROIs) using non-generative sampling techniques as demonstrated in 13. In comparison to these prior works, ours is the first to demonstrate driving at an explanatory level, not only identifying stimuli that can drive a particular voxel or ROI, but also hypothesizing the shared semantic features that cause the underlying activity. The methodology here runs parallel to methods for causal interventions in mechanistic interpretability [41, 42, 43], which have been applied to understanding artificial neural networks.
Our work can be applied to generating a new class of stimuli for online, causal experiments of language processing. We consider this work a piece of a larger online feedback system that continually refines its understanding of brain selectivity by generating hypotheses, testing these hypotheses with new experiments, and refining the hypotheses based on the experimental results. As an example, such a framework could be an especially effective tool in lesion case studies where general observations must be made efficiently from relatively little data.
While effective, GEM-V has several limitations. Most notably, it focuses on a single explanation, missing voxel activity that is polysemantic, i.e. that is not driven well by any single explanation. See Section A.2 for a more detailed discussion of this limitation and how future work might overcome it. Second, the effectiveness of GEM-V depends on staying within the manifold of stimuli trained on during encoding model training. This limitation is explored further in A.6 where we show that poorly generated, off-manifold stimuli are more difficult to predict and yield inferior driving outcomes. GEM-V is also sensitive to the properties of the underlying encoding models–using different encoding models could yield different explanations which might be equally valid. Thus we must be cautious not to interpret the explanations given by GEM-V as uniquely capturing the functional properties of a voxel or ROI. However, this issue will become less severe over time as better and more diverse encoding models become available.
The work here opens the door to a great deal of future research. One avenue should explore whether the degree of granularity of the explanation that GEM-V generates is appropriate for cortical language selectivity. For example, one of the explanations proposed by GEM-V is birthdays, but it would be somewhat surprising to have a specifically birthday-selective area in human cortex. Rather, it is more plausible that such a voxel is selective for something more generic, such as social gatherings [44]. Such ambiguities are of the type that are liable to be resolved by higher-fidelity modelling or iterative causal experiments, such as the hypothesis-refinement feedback system proposed above. GEM-V is also fundamentally limited by the spatiotemporal granularity of the scanning modality we test it on, fMRI. Applying GEM-V with faster or more fine-grained scanning modalities, such as 7T laminar fMRI, ECoG, or NeuroPixels, could reveal selectivity patterns that cannot be effectively observed through fMRI.
GEM-V demonstrates a general framework for aligning the data-driven predictions of models to the theory-oriented hypotheses of classical science. In this paper, we show that applying techniques from modern LLM interpretability research can improve our understanding of cognitive brain processes. Through simple prompting-based methods, we successfully recover semantic selectivity hypotheses that have previously been the providence of painstaking controlled experiments created over decades, and open the door to many more similar results. Our framework is extensible to other modalities such as vision and - requiring only naturalistic data from a single subject - can be easily adapted for use in case study analyses. Furthermore, GEM-V is data-driven and requires minimal human input, countering a frequent criticism of such theories as cherry-picked or driven by preexisting scientific biases. GEM-V will help move toward a future with increasingly nuanced, operationalized, and falsifiable explanations in language neuroscience.
8   Methods
Data collection Two sets of MRI data are used in this study: the first for fitting encoding models and the second for selectively driving brain regions. The original set is described and made openly available in previous work [11, 45]; we describe the details for the second newly collected set here. functional magnetic resonance imaging (fMRI) data were collected from the same 3 human subjects as the original set as they listened to English-language podcast stories over Sensimetrics S14 headphones. Subjects were not asked to make any responses, but simply to listen attentively to the stories. All subjects were healthy and had normal hearing. The experimental protocol was approved by the Institutional Review Board at the University of Texas at Austin. Written informed consent was obtained from all subjects.
All MRI data were collected on a 3T Siemens Skyra scanner at University of Texas at Austin using a 64-channel Siemens volume coil. Functional scans were collected using a gradient echo EPI sequence with repetition time (TR) = 2.00 s, echo time (TE) = 30.8 ms, flip angle = 71°, multi-band factor (simultaneous multi-slice) = 2, voxel size = 2.6mm x 2.6mm x 2.6mm (slice thickness = 2.6mm), matrix size = 84x84, and field of view = 220 mm. Anatomical data were collected using a T1-weighted multi-echo MP-RAGE sequence with voxel size = 1mm x 1mm x 1mm following the Freesurfer morphometry protocol  [46].
Data preprocessing All functional data were motion corrected using the FMRIB Linear Image Registration Tool (FLIRT) from FSL 5.0. FLIRT was used to align all data to a template that was made from the average across the first functional run in the first story session for each subject. These automatic alignments were manually checked for accuracy.
Low frequency voxel response drift was identified using a 2nd order Savitzky-Golay filter with a 120 second window and then subtracted from the signal. To avoid onset artifacts and poor detrending performance near each end of the scan, responses were trimmed by removing 20 seconds (10 volumes) at the beginning and end of each scan, which removed the 10-second silent period and the first and last 10 seconds of each story. The mean response for each voxel was subtracted and the remaining response was scaled to have unit variance.
Voxelwise encoding models Encoding models were fit for each of the three subjects on listening data from roughly 20 hours of unique stories across 20 scanning sessions, yielding a total of $\sim$33,000 datapoints for each voxel across the whole brain. For model testing, the subjects listened to the two test stories five times each, and one test story 10 times, at a rate of 1 test story per session. These test responses were averaged across repetitions.
For each subject and each voxel, we fit a separate encoding model to predict the BOLD response from the features we extract from the stimulus. Features were extracted from the 18th layer of the 30-billion parameter LLaMA model [13], and the 33rd layer of the 30-billion parameter OPT model [12]. These layers were chosen based on prior work [8] that determined that they were the most performant at predicting brain activity. To temporally align word times with TR times, Lanczos interpolation was applied with a window size of 3. The hemodynamic response function was approximated with a finite impulse response model using 4 delays at -8,-6,-4 and -2 seconds [20].
To evaluate the voxelwise encoding models, we used the learned encoding model to generate and evaluate predictions on a held-out test set. The OPT features achieved a mean voxelwise correlation of about 0.128 whereas the LLaMA features achieved a mean voxelwise correlation of about 0.132. These performances exceed that of previously published models on the same dataset (mean correlation 0.111) that were able to produce meaningful semantic decoding [45].
Voxel selection for explanation We selected 500 well-modeled, diverse voxels to explain for each subject. To ensure that these voxels were well-modeled, we selected only from voxels with a test correlation above 0.15, (this corresponds to the top 4̃0% most well-predicted voxels). Then, we applied principal components analysis to the learned ridge weights across all well-modeled voxels in cortex. We project all voxels to the first four principal components, which are known to encode differences in semantic selectivity [20] and then uniformly sampled voxels from the within the convex hull in $\mathcal{R}^{4}$ induced by this projection. This uniform sampling ensures diversity in the selected voxels. The mean voxel correlation for the 1,500 voxels we study is 0.35.
Framework for generating explanations GEM-V yields a short, natural-language explanation describing what elicits the strongest response from each of the 1,500 selected voxel encoding models. To obtain this explanation, we follow the summarize and score (SASC) framework introduced in previous work [17] (see Fig. 1b). SASC first generates candidate explanations (using GPT-4 [16]) based on the n-grams ($n=1,2,3$) that elicit the most positive response from the LLaMA encoding model. Each candidate explanation is then evaluated by generating synthetic data based on the explanation and testing the response of the encoding model to the data.
Stability score screening After extracting an explanation for 500 voxels per subject, we selected 17 voxels per subject for followup fMRI experiments. 17 voxels were selected so that the resulting 17-paragraph stories would be similar in length to the stimuli used to initially fit the encoding models [11]. To select voxels which are not only well-predicted but also likely to be well-explained, we introduced the stability score, which measures the correlation of predictions by the encoding models based on OPT and LLaMA for each unique n-gram present in the dataset of stories. A higher correlation implies greater agreement between the different encoding models and thus greater stability. For each subject, we first filter the 40 voxels with the highest stability score and then manually select 17 voxels from this set that have diverse explanations.
Story generation (single-voxel setting) For follow-up experiments, we prompted a large LLM (GPT-4 [16]) to generate stories based on the explanations for our selected voxels. Stories are generated by repeatedly prompting the LLM in a chat to continue the story, one paragraph at a time. For each paragraph, the LLM is asked to focus on one explanation and to include related key n-grams (Fig. 4a, see full prompts in Section A.1).
Before running follow-up experiments, we checked that each paragraph’s text matched its generated explanation (Fig. 4a). This match was measured by prompting an LLM [47] to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation. We also validated that each story paragraph drives the encoding model for its corresponding voxel (Fig. 4b). We generated 8 different stories by changing the random seed for each subject and kept the best 2 for S01, the best 6 for S02 (the pilot subject), and the best 2 for S03.
ROI setting The ROI setting uses the same framework as the single-voxel setting, but instead of maximizing the response in a single voxel, we sought to maximize the average response of the encoding model outputs over all voxels in an ROI. When selecting explanations, we used the average outputs of the OPT and LLaMA encoding models to help make results more stable.
Appendix A Extended Data
A.1 Prompting details
Prompts used for explanation
Following the original SASC study [17], we use the following prompts for explanation.
[…]
In the synthetic scoring step, we generate similar synthetic strings with the prompt Generate 10 phrases that are similar to the concept of {explanation}:. For dissimilar synthetic strings we use the prompt Generate 10 phrases that are not similar to the concept of {explanation}:. Minor automatic processing is applied to LLM outputs, e.g. parsing a bulleted list, converting to lowercase, and removing extra whitespaces.
Prompts used for story generation
We use two variations of prompts. In the main setting, we begin with Write the beginning paragraph of a long, coherent story. The story should be about ”expl”. Make sure it contains several words related to ”expl”, such as examples. Between paragraphs, it changes to Write the next paragraph of the story, staying consistent with the story so far, but now make it about…
In the second setting, we instead begin with Write the beginning paragraph of an interesting story told in first person. The story should have a plot and characters. The story should be about… and continue with Write the next paragraph of the story, but now make it about…
Prompts used for ROI-driving story generation
For driving ROIs (Section 4), many explanations are related to locations, and so we append a suffix to the prompt for reach paragraph (see Table 1) to help make paragraphs more distinct.
Table 1: Suffixes added to the prompt for generating the story paragraph for each ROI.
ROI Driving explanation Suffix
IPS Descriptive elements of scenes or objects Avoid mentioning any locations.
OFA Conversational transitions Avoid mentioning any locations.
OPA Direction and location descriptions Avoid mentioning any specific location names (like ”New York” or ”Europe”).
OPA-only Self-reflection and growth Avoid mentioning any specific location names (like ”New York” or ”Europe”).
PPA Scenes and settings Avoid mentioning any specific location names (like ”New York” or ”Europe”).
PPA-only Unappetizing foods Avoid mentioning any specific location names (like ”New York” or ”Europe”).
RSC Travel and location names
RSC-only Location names
sPMv Dialogue and responses Avoid mentioning any locations.
A.2 Driving polysemantic voxels
We additionally run experiments seeking to drive multiple descriptions corresponding to the same voxel. These experiments generate and test GEM-V stories in the same manner as testing individual explanations (Fig. 1c) but differ in the way they source voxel descriptions. In our first investigation (with subject S02), we select two explanations that both come from SASC, in the case that its summaries of the top ngrams yield two distinct explanations. In our second investigation (now with subject S03), we instead select two explanations that each come from a different encoding model. In both cases, each story tests 8 voxels with two explanations each. We average the results over two stories per subject.
We generally find that we are able to only successfully drive one of the two explanations we seek to drive for an individual voxel (Fig. 5). We believe this is due to SASC’s overreliance on summarizing top-driving ngrams, and could potentially be mitigated by summarizing a broader range of encoding model inputs.
Figure 5: Results for driving polysemantic voxels. (a) Voxel response for driving paragraphs (blue) show a small increase relative to the baseline responses of the remaining paragraphs (gray). Each voxel appears as two points connected by a vertical line corresponding; each point shows the result when driving the voxel using a different explanation. Most voxels are only driven successfully for a single explanation.
A.3 Story-level breakdowns
In this section, we show results for GEM-V at the level of individual stories. Fig. 6 shows the mean driving voxel response across all subjects and settings. Fig. 7 stratifies these synthetic stories based on the underlying prompt used to generate the stories and Fig. 8 shows results for these stories at driving voxel clusters rather than individual voxels.
Figure 6: Different stories vary in their ability to drive voxels. There is large variance between the driving response produced for different stories, even for the same subject. Driving response shows the difference from baseline paragraphs. For ROI stories, this shows the average response of voxels in an ROI; for other stories this shows the average response of the single target voxel.
Figure 7: Different prompts yield different driving performance. The driving ability of the generated story differs based on the precise wording of the prompt. We compare two versions of the prompt, and find that results for version 1 are noticeably better. Version 1 is our default prompt, described in Section A.1. Version 0 instead begins with Write the beginning paragraph of an interesting story told in first person. The story should have a plot and characters. The story should be about.. and transitions between paragraphs with the prompt Write the next paragraph of the story, but now make it about…. Results show the 6 stories run for S02 in the default setting.
Figure 8: Clusters of voxels semantically related to the voxel cluster are driven, but less reliably than the target voxels themselves. Voxel clusters group semantically similar voxels by using the learned ridge regression weights of the encoding model. Results show the 6 stories run for S02 in the default setting.
Table 2: Head motion for different subjects during GEM-V stories. S01 shows substantially larger motion than the other two subjects.
Subject Average framewise displacement (mm) Fraction of TRs with framewise displacement above 0.5 mm
S01 0.273 0.094
S02 0.108 0.000
S03 0.102 0.000
A.4 Correlates of driving performance at the voxel level
In this section we provide additional analyses of what voxel-level factors inform whether driving performance will succeed. The strongest correlate is the stability score, given in the main text (Fig. 4c).
a b
Refer to caption Refer to caption
Figure 9: Correlates of driving performance at the voxel level. (a) Ngram matching score, i.e. how well the driving paragraphs match their generating explanations, does not correlate well with driving score. (b) Encoding driving score, i.e. how well the driving paragraphs succeed in driving the encoding model, correlates well with driving score for 2 of the 3 subjects.
S01 S03
Refer to caption Refer to caption
Figure 10: To confirm that generated paragraphs match the explanation used to construct them, a matching score was computed for each explanation and paragraph by using an LLM to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation and then z-scoring the result. Each driving paragraph showed a strong match with its generating explanation. Plot shows subject S01 and S03; plot for subject S02 shown in Fig. 4a.
        S01              S03
Refer to caption Refer to caption
Figure 11: To confirm that each driving paragraph effectively drives its corresponding encoding model, we computed the predicted response in each selected voxel to each generated paragraph. This revealed strong matches for most voxels, along with some matches between driving paragraphs and voxels with semantically similar explanations, e.g. directions and locations. Plot shows subject S01 and S03, plot for subject S02 shown in Fig. 4b.
A.5 Selectively driving pairs of voxels
Useful and explanatory scientific theories are robust to unrelated distributional shifts. For example, if a region or voxel is selective for food concepts, then that region should remain selective for food concepts even if those concepts cooccur with unrelated semantic concepts. In order to test this independence of our generated explanations, we test whether GEM-V can be used to simultaneously drive pairs of voxels in different areas across the brain.
Fig. 12a shows the the multi-voxel driving pipeline. Pairs of voxels with semantically independent generated explanations are selected and stimuli are generated so that the stimuli have the following pattern: (i) the first voxel is driven without respect to the second voxel by using the explanation for that voxel, (ii) then both are driven simultaneously with a generated stimulus that contains both generated explanations, and then (iii) only the second voxel is driven without respect to the first voxel. Fig. 12b demonstrates that we are able to effectively combine the voxel-level explanations to drive separate voxels at the same time, or independently drive one without driving the other. This shows that the generated explanations are resilient to unrelated semantic interventions, demonstrating that they do not work merely because they tend to semantically co-occur with stimuli that actually drives those voxels.
Figure 12: LLM stories can selectively drive fMRI responses for pairs of voxels. (a) The pipeline for selectively driving a pair voxels generates explanations on a per voxel basis and prompts an LLM to generate paragraphs that either drive an individual voxel or a pair of voxels. (b) Voxel response when alternating between driving a single voxel (blue) or a pair of voxels (purple) generally succeeds in driving relative to the baseline (gray dotted line).
A.6 Checkerboard driving
Figure 13: (a) A suitable location for a target checkerboard pattern is chosen from prefrontal cortex by cross-referencing the target pattern with a pre-generated fMRI passive listening dataset. (b) A subject listens to a story that elicits high variance for the checkerboard pattern. We reconstruct the target pattern from the data by weighting the response at each timepoint by its predicted similarity to the target pattern. The reconstruction partially recovers the target pattern, achieving reasonably high correlation with the target pattern.
To test the limits of our voxel-driving paradigm, we further examine whether they can be used directly to produce arbitrary activation patterns if explanatory considerations were dropped. The experimental pipeline for this test is shown in Fig. 13a. We select a location on the cortical surface to drive by searching our training dataset for areas that contain at least one response from our training dataset with a high cosine similarity to a target checkerboard pattern. We choose a checkerboard pattern for this experiment to demonstrate that driving can be achieved even for relatively complex targets, so long as the target is biologically possible.
We demonstrate that we can partially manifest a checkerboard pattern in prefrontal cortex. As the checkerboard pattern is fairly unnatural, we aim to reconstruct the checkerboard from the response pattern rather than directly drive it (Fig. 13b). We compute the sum of the responses for the voxels in the checkerboard location across a single high-variance story, weighted by the cosine similarity of the responses to the target pattern. This results in a pattern that bears some resemblance to the desired checkerboard (pearson correlation coefficient $\rho=0.67$).
A.7 Semantic explanation average constrast maps [^1] Full set of flatmaps can also be found at github.com/microsoft/automated-explanations
Figure 14: Explanation contrast for all tested explanations in S01. Arrows indicate source voxel where driving was attempted.
Figure 15: Explanation contrast for all tested explanations in S02. Arrows indicate source voxel where driving was attempted.
Figure 16: Explanation contrast for all tested explanations in S03. Arrows indicate source voxel where driving was attempted.